OpenStudy (anonymous):

use the method of elimination to solve the system of linear equations. x+y=3 x-y=-7

hero (hero):

y = 3-x y = x+7 y = y 3-x = x+7 2x = -4 x = -2 y = 5 The two linear equations intersect geometrically at the point (-2,5)